Circular No. 6087

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994Z IN NGC 87 A. Wassilieff, Palmerston North, New Zealand, reports his CCD dis-

covery of a supernova on Oct. 2.51 UT, ∼ 36′′ east and ∼ 12′′ south of the nucleus of NGC 87. The object was estimated at mag 14.6 and noted again on Oct. 3.38, and it was not present on a similar image (scale 7′′.5/mm) obtained on 1992 Nov. 29. Measurements by G. V. Williams of the faxed image yield the position α = 0^{h}18^{m}49^{s}.49, δ = −48^{o}54′28′′.0 (equinox 1950.0). End figures for the galaxy nucleus are 47^{s}.16, 20′′.3.

M. M. Phillips, Cerro Tololo Interamerican Observatory, reports that CCD images obtained by J. Maza and X. Gomez on Oct. 4 confirm the presence of an apparent supernova in NGC 87. Conditions were not photometric, but a comparison with a sequence around T Phe gave V ∼ 16, B − V ∼ +0.1.

S. Benetti, European Southern Observatory (ESO) and G. Hasinger, Astrophysikalisches Institut, Potsdam, report that inspection of a fullyreduced CCD spectrum (range 450–700 nm, resolution 1.1 nm) obtained on Oct. 4.12 UT with ESO 2.2-m telescope + EFOSC2 confirms that the object is a type II supernova 2–3 weeks after outburst. Strong Hβ and Hα lines with P-Cyg profiles are superimposed on a blue continuum. The He I 587.6-nm line is also present with a P-Cyg profile. The expansion velocity deduced from the minima of the Hβ and He I 587.6-nm lines is ∼ 10 000 km/s. A velocity of 12 000 km/s is deduced from the Hα minimum.

X-RAY NOVA IN SCORPIUS N. Alexandrovich, K. Borozdin, V. Efremov and R. Sunyaev, Space

Research Institute, Moscow, on behalf of the MIR-KVANT team, report: “According to data from the TTM telescope onboard the MIR space station this x-ray nova is in a high (soft) spectral state typical of black-hole candidates and similar to that observed recently in the superluminous radio source GRS 1915−105 (IAUC 6080). The flux detected at 5 keV was 500 mCrab on Sept. 24 and 600 mCrab on Sept. 25. The source spectrum within the 2–15-keV band can be approximated by a blackbody fit with kT = 1.06 ± 0.02 keV. According to MIR-KVANT-HEXE data, the 3σ upper limit to the tail of the hard flux is 5 and 15 mCrab in the 15–45keV and the 45–105-keV data, respectively. In hard x-rays the source is much weaker than 4U 1700−37 (50 mCrab in 15–45-keV band), which is separated from the Sco x-ray nova by < 3^{o}.”

1994 October 4 Brian G. Marsden